A website had 120 visitors on its first day.The number of visitors double in every 4 days.Write a function comparing the number of visitors for the particular number of days.
Answer by stanbon(75887)

A website had 120 visitors on its first day.The number of visitors double in every 4 days.Write a function comparing the number of visitors for the particular number of days.
----
Form:: y = ab^x
------
y = 120(2)^(d/4)
----
f(0) = 120
f(1) = 120*2^(1/4) = 143 when rounded up
f(2) = 120*2^(2/4) = 170 when rounded up
f(3) = 120*2^(3/4) = 202 when rounded up
f(4) = 120*2^(4/4) = 240
--------------
Cheers,
Stan H.
--------------